How do you solve the missing numbers of the two multiplication problems?
1). 10 x ________ = 4,276.4
2). 0.1 x ________ = 7,354

Answer by Theo(13342)

How do you solve the missing numbers of the two multiplication problems?

1). 10 x ________ = 4,276.4

if 10 * ____________ = 4,276.4, then 4,276.4 / 10 = _____________.
solve for ______________ to get ___________ = 4,276.4 / 10 = 427.64.

2). 0.1 x ________ = 7,354

if 0.1 * _____________ = 7,354, then 7,354 / 0.1 = __________________.
solve for __________________ to get _______________ = 7,354 / 0.1 = 73,540.

you can use your calculator to confirm.

4,276.4 / 10 = 427.64.
7,354 / 0.1 = 73,540.